the gate receipts at the basketball game in gym is 43,500 pesos for 2 500 spectators. If the price for reserved seat was 20 pesos and the general admission was 15 pesos, how many ticket were sold each kind?
Answer by Maths68(1474)

Total amount for selling tickets for general and reserved seats = 43,500 pesos
price of a ticket for reserved seat = 20 pesos
price of a ticket for general admission = 15 pesos,
Total = 2500 spectators or tickets
Let
Tickets sold for reserved seat = x
Tickets sold for general admission = y
20x+15y=43500..............(1)
x+y=2500........................(2)
x+y=2500
x=2500-y........................(3)
Put the value of x in (1)
20(2500-y )+15y=43500
50000-20y+15y=43500
50000-5y=43500
-5y=43500-50000
-5y=-6500
Multiply by (-1) both sides
5y=6500
Divid by 5 both sides
5y/5=6500/5
Y=1300
Put the value of y in (3)
x=2500-y
x=2500-1300
x=1200
Tickets sold for reserved seat=1200
And
Tickets sold for general admission=1300